For the experiment of drawing a single card from a standard 52-card deck, find (a) the probability, and (b) the odds in favor, of each event.

(1) Not an ace
(2) Club or heart
(3) Not a heart or a seven

Answer by stanbon(75887)

For the experiment of drawing a single card from a standard 52-card deck, find (a) the probability, and (b) the odds in favor, of each event.
I'll do the 1st two:
----
(1) Not an ace
P(no ace) = 48/52;P(ace) = 4/52
Odds in favor of = [P(no ace)]/[P(ace)] = 48/4 = 12:1
------------------------------------------
(2) Club or heart
P(c or h) = P(c) + P(h) = 1/4 + 1/4 = 1/2
P(no c or h) = 1/2
odds in favor: = P(c or h)/P(no c or h) = 1/2 / 1/2 = 1:1
------------------------------------------
(3) Not a heart or a seven
===============================
Cheers,
Stan H.
===============================